### v4

# The computational power of discrete chemical reaction networks with bounded executions

Abstract
Chemical reaction networks (CRNs) model systems where molecules interact according to a finite set of reactions such as $A+B\to C$, representing that if a molecule of $A$ and $B$ collide, they disappear and a molecule of $C$ is produced. CRNs can compute Boolean-valued predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ and integer-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}$; for instance $X_{1}+X_{2}\to Y$ computes the function $\min(x_{1},x_{2}),$ since starting with $x_{i}$ copies of $X_{i}$, eventually $\min(x_{1},x_{2})$ copies of $Y$ are produced.
We study the computational power of execution bounded CRNs, in which only a finite number of reactions can occur from the initial configuration (e.g., ruling out reversible reactions such as $A\mathop{\rightleftharpoons}\limits B$). The power and composability of such CRNs depends crucially on some other modeling choices that do not affect the computational power of CRNs with unbounded executions, namely whether an initial leader is present, and whether (for predicates) all species are required to “vote” for the Boolean output. If the CRN starts with an initial leader, and can allow only the leader to vote, then all semilinear predicates and functions can be stably computed in $O(n\log n)$ parallel time by execution bounded CRNs.
However, if no initial leader is allowed, all species vote, and the CRN is “non-collapsing” (does not shrink from initially large to final $O(1)$ size configurations), then execution bounded CRNs are severely limited, able to compute only eventually constant predicates. A key tool is a characterization of execution bounded CRNs as precisely those with a nonnegative linear potential function that is strictly decreased by every reaction [czerner2024fast].
keywords:
chemical reaction networks, population protocols, stable computation
1 Introduction
Chemical reaction networks (CRNs) are a fundamental tool for understanding and designing molecular systems. By abstracting chemical reactions into a set of finite, rule-based transformations, CRNs allow us to model the behavior of complex chemical systems. For instance, the CRN with a single reaction $2X\rightarrow Y$, produces one $Y$ every time two $X$ molecules randomly react together, effectively calculating the function $f(x)=\lfloor x/2\rfloor$ if the initial count of $X$ is interpreted as the input and the eventual count of $Y$ as the output. A commonly studied special case of CRNs is the population protocol model of distributed computing [angluin2007computational], in which each reaction has exactly two reactants and two products, e.g., $A+B\to C+D$. This model assumes idealized conditions where reactions can proceed indefinitely, constrained only by the availability of reactants in the well-mixed solution.
Precisely the semilinear predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ [angluin2004computation] and functions $f:\mathbb{N}^{d}\to\mathbb{N}$ [Chen2012DeterministicFunction] can be computed stably, roughly meaning that the output is correct no matter the order in which reactions happen. In population protocols or other CRNs with a finite reachable configuration space, this means that the output is correct with probability 1 under a stochastic scheduler that picks the next molecules to react at random. However, existing constructions to compute semilinear predicates and functions use CRNs with unbounded executions, meaning that it is possible to execute infinitely many reactions from the initial configuration. CRNs with bounded executions have several advantages. With an absolute guarantee on how many reactions will happen before the CRN terminates, wet-lab implementations need only supply a bounded amount of fuel to power the reactions. Such CRNs are simpler to reason about: each reaction brings it “closer” to the answer. They also lead to a simpler definition of stable computation than is typically employed: an execution bounded CRN stably computes a predicate/function if it gets the correct answer after sufficiently many reactions.
To study this topic, we study networks that must eventually reach a configuration where no further reactions can occur, regardless of the sequence of reactions executed. This restriction is nontrivial because the techniques of [Chen2012DeterministicFunction, Doty2013Leaderless] rely on reversible reactions (leading to unbounded executions) catalyzed by species we expect to be depleted once a computational step has terminated. This trick seems to add computational power to our system by undoing certain reactions as long as a specific species is present. Consider the following CRN computing $f(x_{1},x_{2},x_{3})=\min\left(x_{1}-x_{2},x_{3}\right)$. The input values $x_{i}$ are given by the counts of $X_{i}$, and the output by the count of $Z$ molecules in the stable state:
$X_{1}$ $\rightarrow Y$ (1)
$X_{2}+Y$ $\rightarrow\varnothing$ (2)
$Y+X_{3}$ $\rightarrow Z$ (3)
$Z+X_{2}$ $\rightarrow X_{2}+X_{3}+Y$ (4)
Reactions (1) and (2) compute $x_{1}-x_{2}$, storing the result in the count of $Y$. Next, reaction (3) can be applied exactly $\min(y,x_{3})$ times. But since the order of reactions is a stochastic process, we might consume copies of $Y$ in (3), before all of $x_{2}$ is subtracted from it. Therefore, we add reaction (4), using $X_{2}$ as a catalyst to undo reaction (3) as long as copies of $X_{2}$ are present, indicating that the first step of computation has not terminated. However, this means the above CRN does not have bounded exectutions, since reactions (3) and (4) can be alternated in an infinite execution. A similar technique is used in [Chen2012DeterministicFunction], where semilinear sets are understood as a finite union of linear sets, shown to be computable in parallel by CRNs. A reversible, catalyzed reaction finally converts the output of one of the CRNs to the global output. Among other questions, we explore how the constructions of [Chen2012DeterministicFunction] and [Doty2013Leaderless] can be modified to provide equal computational power while guaranteeing bounded execution.
The paper is organized as follows. Section˜3 defines execution boundedness (Definition˜3.1). We introduce alternative characterizations of the class for use in later proofs, such as the lack of self-covering execution paths. Section˜4 and 5 contain the main positive results of the paper and provide the concrete constructions used to decide semilinear sets and functions using execution bounded CRNs whose initial configurations contain a single leader. Section˜6 discusses the limitations of execution bounded CRNs, introducing the concept of a “linear potential function” as a core characterization of these systems. We demonstrate that entirely execution bounded CRNs that are leaderless and non-collapsing (such as all population protocols), can only stably decide trivial semilinear predicates: the eventually constant predicates (Definition˜6.9).
2 Preliminaries
We use established notation from [Chen2012DeterministicFunction, Doty2013Leaderless] and stable computation definitions from [angluin2007computational] for (discrete) chemical reaction networks.
2.1 Notation
Let $\mathbb{N}$ denote the nonnegative integers. For any finite set $\Lambda$, we write $\mathbb{N}^{\Lambda}$ to mean the set of functions $f:\Lambda\rightarrow\mathbb{N}$. Equivalently, $\mathbb{N}^{\Lambda}$ can be interpreted as the set of vectors indexed by the elements of $\Lambda$, and so $\mathbf{c}\in\mathbb{N}^{\Lambda}$ specifies nonnegative integer counts for all elements of $\Lambda$. $\mathbf{c}(i)$ denotes the $i$-th coordinate of $\mathbf{c}$, and if $\mathbf{c}$ is indexed by elements of $\Lambda$, then $\mathbf{c}(Y)$ denotes the count of species $Y\in\Lambda$. We sometimes use multiset notation for such vectors, e.g., $\{A,3C\}$ for the vector $(1,0,3)$, assuming there are three species $A,B,C$. If $\Sigma\subseteq\Lambda$, then $\mathbf{i}\upharpoonright\Sigma$ denotes restriction of $\mathbf{i}$ to $\Sigma$.
[…]
For a matrix or vector $\mathbf{x}$, define $\|\mathbf{x}\|=\|\mathbf{x}\|_{1}=\sum_{i}|\mathbf{x}(i)|$, $i$ ranges over all the entries of $\mathbf{x}$.
2.2 Chemical Reaction Networks
A chemical reaction network (CRN) is a pair $\mathcal{C}=(\Lambda,R)$, where $\Lambda$ is a finite set of chemical species, and $R$ is a finite set of reactions over $\Lambda$, where each reaction is a pair $(\mathbf{r},\mathbf{p})\in\mathbb{N}^{\Lambda}\times\mathbb{N}^{\Lambda}$ indicating the reactants $\mathbf{r}$ and products $\mathbf{p}$. A population protocol [angluin2004computation] is a CRN in which all reactions $(\mathbf{r},\mathbf{p})$ obey $\|\mathbf{r}\|=\|\mathbf{p}\|=2$. (Note that CRNs, including population protocols, do not assume any underlying “communication graph” and model a well-mixed system in which each equal-sized of molecules is as likely to collide and react as any other.) We write reactions such as $A+2B\mathop{\rightarrow}\limits A+3C$ to represent the reaction $(\{A,2B\},\{A,3C\})$. A configuration $\mathbf{c}\in\mathbb{N}^{\Lambda}$ of a CRN assigns integer counts to every species $S\in\Lambda$. When convenient, we use the notation $\left\{n_{1}S_{1},n_{2}S_{2},\ldots,n_{k}S_{k}\right\}$ to describe a configuration $\mathbf{c}$ with $n_{i}\in\mathbb{N}$ copies of species $S_{i}$, i.e., $\mathbf{c}(S_{i})=n_{i}$, and any species that is not listed is assumed to have a zero count. If some configuration $\mathbf{c}$ is understood from context, for a species $S$, we write #S\#S to denote $\mathbf{c}(S).$ A reaction $(\mathbf{r},\mathbf{p})$ is said to be applicable in configuration $\mathbf{c}$ if $\mathbf{r}\leqq\mathbf{c}$. If the reaction $(\mathbf{r},\mathbf{p})$ is applicable, applying it results in configuration $\mathbf{c}^{\prime}=\mathbf{c}-\mathbf{r}+\mathbf{p}$, and we write $\mathbf{c}\rightarrow\mathbf{c}^{\prime}$.
An execution $\mathcal{E}$ is a finite or infinite sequence of one or more configurations $\mathcal{E}=(\mathbf{c}_{0},\mathbf{c}_{1},\mathbf{c}_{2},\ldots)$ such that, for all $i\in\{1,\ldots,|\mathcal{E}|-1\},\mathbf{c}_{i-1}\rightarrow\mathbf{c}_{i}$ and $\mathbf{c}_{i-1}\neq\mathbf{c}_{i}$. $\mathbf{x}\Rightarrow_{P}\mathbf{y}$ denotes that $P$ is finite, starts at $\mathbf{x}$, and ends at $\mathbf{y}$. In this case we say $\mathbf{y}$ is reachable from $\mathbf{x}$. Let $\mathsf{reach}(\mathbf{x})=\{\mathbf{y}\mid\mathbf{x}\Rightarrow\mathbf{y}\}$. Note that the reachability relation is additive: if $\mathbf{x}\Rightarrow\mathbf{y}$, then for all $\mathbf{c}\in\mathbb{N}^{\Lambda}$, $\mathbf{x}+\mathbf{c}\Rightarrow\mathbf{y}+\mathbf{c}$.
For a CRN $\mathcal{C}=(\Lambda,R)$ where $|\Lambda|=n$ and $|R|=m$, define the $n\times m$ stoichiometric matrix $\mathbf{M}$ of $\mathcal{C}$ as follows. The species are ordered $S_{1},\dots,S_{n}$, and the reactions are ordered $(\mathbf{r}_{1},\mathbf{p}_{1}),\dots,(\mathbf{r}_{m},\mathbf{p}_{m})$, and $\mathbf{M}_{ij}=\mathbf{p}_{j}(S_{i})-\mathbf{r}_{j}(S_{i})$. In other words, $\mathbf{M}_{ij}$ is the net amount of $S_{i}$ produced when executing the $j$’th reaction. For instance, if the CRN has two reactions $S_{1}\mathop{\rightarrow}\limits S_{2}+2S_{3}$ and $3S_{2}+S_{3}\mathop{\rightarrow}\limits S_{1}+S_{2}+S_{3}$, then
$\mathbf{M}=\begin{pmatrix}-1&1\\ 1&-2\\ 2&0\end{pmatrix}$.
Remark 2.1.
Let $\mathbf{u}\in\mathbb{N}^{R}$. Then the vector $\mathbf{M}\mathbf{u}\in\mathbb{Z}^{\Lambda}$ represents the change in species counts that results from applying reactions by amounts described in $\mathbf{u}$. In the above example, if $\mathbf{u}=(2,1)$, then $\mathbf{M}\mathbf{u}=(-1,0,4)$, meaning that executing the first reaction twice ($\mathbf{u}(1)=2$) and the second reaction once ($\mathbf{u}(2)=1$) causes $S_{1}$ to decrease by 1, $S_{2}$ to stay the same, and $S_{3}$ to increase by 4.
2.3 Stable computation with CRNs
To capture the result of computations done by a CRN, we generalize the definitions to include information about how to interpret the final configuration after letting the CRN run until the result cannot change anymore (characterized below as stable computation). Computation primarily involves two classes of functions: 1. evaluating predicates $\phi:\mathbb{N}^{k}\to\{0,1\}$ to determine properties of the input, and 2. executing general functions that map an input configuration to an output, denoted as $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$.
The definitions below reference input species $\Sigma\subseteq\Lambda$ and an initial context $\mathbf{s}\in\mathbb{N}^{\Lambda\setminus\Sigma}$. If $\mathbf{s}=\mathbf{0}$ we say that CRN is leaderless. The initial context may be any constant multiset of species, though in practice it tends to be a single “leader” molecule. Furthermore, other initial contexts such as $\{2A,3B\}$ could be produced from a single leader $L$ via a reaction $L\to 2A+3B$, so we may assume without loss of generality that the initial context, if it is nonzero, is simply a single leader. In both cases, we say $\mathbf{i}\in\mathbb{N}^{\Lambda}$ is a valid initial configuration if $\mathbf{i}=\mathbf{s}+\mathbf{x}$, where $\mathbf{x}(S)=0$ for all $S\in\Lambda\setminus\Sigma$; i.e., $\mathbf{i}$ is the initial context plus only input species.
A chemical reaction decider (CRD) is a tuple $\mathcal{D}=(\Lambda,R,\Sigma,\Upsilon_{1},\Upsilon_{0},\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subseteq\Lambda$ is the set of input species, $\Upsilon_{1}\subseteq\Lambda$ is the set of yes voters, and $\Upsilon_{0}\subseteq\Lambda$ is the set of no voters, such that $\Upsilon_{1}\cap\Upsilon_{0}=\emptyset$, and $\mathbf{s}\in\mathbb{N}^{\Lambda\backslash\Sigma}$ is the initial context. If $\Upsilon_{1}\cup\Upsilon_{0}=\Lambda$, we say the CRD is all-voting. We define a global output partial function $\Phi:\mathbb{N}^{\Lambda}\dashrightarrow\{0,1\}$ as follows. $\Phi(\mathbf{c})$ is undefined if either $\mathbf{c}=\mathbf{0}$, or if there exist $S_{0}\in\Upsilon_{0}$ and $S_{1}\in\Upsilon_{1}$ such that $\mathbf{c}\left(S_{0}\right)>0$ and $\mathbf{c}\left(S_{1}\right)>0$. In other words, we require a unanimous vote as our output. We say $\mathbf{c}$ is stable if, for all $\mathbf{c}^{\prime}$ such that $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$, $\Phi(\mathbf{c})=\Phi(\mathbf{c}^{\prime}).$ We say a CRD $\mathcal{D}$ stably decides the predicate $\psi:\mathbb{N}^{\Sigma}\rightarrow\{0,1\}$ if, for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Lambda}$, letting $\mathbf{i}_{0}=\mathbf{i}\upharpoonright\Sigma$, for all configurations $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$ such that $\mathbf{c}^{\prime}$ is stable and $\Phi\left(\mathbf{c}^{\prime}\right)=\psi\left(\mathbf{i}_{0}\right)$. We associate to a predicate $\psi$ the set $A=\psi^{-1}(1)$ of inputs on which $\psi$ outputs 1, so we can equivalently say the CRD stably decides the set $A.$
A chemical reaction computer $(CRC)$ is a tuple $\mathcal{C}=(\Lambda,R,\Sigma,Y,\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subset\Lambda$ is the set of input species, $Y\in\Lambda\backslash\Sigma$ is the output species, and $\mathbf{s}\in\mathbb{N}^{\Lambda\backslash\Sigma}$ is the initial context. A configuration $\mathbf{o}\in\mathbb{N}^{\Lambda}$ is stable if, for every $\mathbf{c}$ such that $\mathbf{o}\Rightarrow\mathbf{c},\mathbf{o}(Y)=\mathbf{c}(Y)$, i.e. the output can never change again. We say that $\mathcal{C}$ stably computes a function $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ if for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Sigma}$ and any $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{o}$ such that $\mathbf{o}$ is stable and $f(\mathbf{i}\upharpoonright\Sigma)=\mathbf{o}(Y)$.
2.4 Time model
The following model of stochastic chemical kinetics is widely used in quantitative biology and other fields dealing with chemical reactions between species present in small counts [Gillespie77]. It ascribes probabilities to execution sequences, and also defines the time of reactions, allowing us to study the computational complexity of the CRN computation in Sections˜4 and 5. If the volume is defined to be the total number of molecules, then the time model is essentially equivalent to the notion of parallel time studied in population protocols [AngluinAE2008Fast]. In this paper, the rate constants of all reactions are $1$, and we define the kinetic model with this assumption. A reaction is unimolecular if it has one reactant and bimolecular if it has two reactants. We use no higher-order reactions in this paper.
The kinetics of a CRN is described by a continuous-time Markov process as follows. Given a fixed volume $v>0$, the propensity of a unimolecular reaction $\alpha:X\to\ldots$ in configuration $\mathbf{c}$ is $\rho(\mathbf{c},\alpha)=\mathbf{c}(X)$. The propensity of a bimolecular reaction $\alpha:X+Y\to\ldots$, where $X\neq Y$, is $\rho(\mathbf{c},\alpha)=\frac{\mathbf{c}(X)\mathbf{c}(Y)}{v}$. The propensity of a bimolecular reaction $\alpha:X+X\to\ldots$ is $\rho(\mathbf{c},\alpha)=\frac{1}{2}\frac{\mathbf{c}(X)(\mathbf{c}(X)-1)}{v}$. The propensity function determines the evolution of the system as follows. The time until the next reaction occurs is an exponential random variable with rate $\rho(\mathbf{c})=\sum_{\alpha\in R}\rho(\mathbf{c},\alpha)$ (note that $\rho(\mathbf{c})=0$ if no reactions are applicable to $\mathbf{c}$). The probability that next reaction will be a particular $\alpha_{\text{next}}$ is $\frac{\rho(\mathbf{c},\alpha_{\text{next}})}{\rho(\mathbf{c})}$.
The kinetic model is based on the physical assumption of well-mixedness that is valid in a dilute solution. Thus, we assume the finite density constraint, which stipulates that a volume required to execute a CRN must be proportional to the maximum molecular count obtained during execution [SolCooWinBru08]. In other words, the total concentration (molecular count per volume) is bounded. This realistically constrains the speed of the computation achievable by CRNs.
For a CRD or CRC stably computing a predicate/function, the stabilization time is the function $t:\mathbb{N}\to\mathbb{N}$ defined for all $n\in\mathbb{N}$ as $t(n)=$ the worst-case expected time to reach from any valid initial configuration of size $n$ to a stable configuration.
2.5 Semilinear sets, predicates, functions
Definition 2.2.
A set $L\subseteq\mathbb{N}^{d}$ is linear if there are vectors $\mathbf{b},\mathbf{p}_{1},\dots,\mathbf{p}_{k}$ such that $L=\{\mathbf{b}+n_{1}\mathbf{p}_{1}+\dots+n_{k}\mathbf{p}_{k}\mid n_{1},\dots,n_{k}\in\mathbb{N}\}$. A set is semilinear if it is a finite union of linear sets. A predicate $\phi:\mathbb{N}^{d}\to\{0,1\}$ is semilinear if the set $\phi^{-1}(1)$ is semilinear. A function $f:\mathbb{N}^{d}\to\mathbb{N}$ is semilinear if its graph $\{(\mathbf{x},y)\in\mathbb{N}^{d+1}\mid f(\mathbf{x})=y\}$ is semilinear.
The following is a known characterization of the computational power of CRNs [angluin2007computational, chen2023rate].
Theorem 2.3 ([angluin2007computational, chen2023rate]).
A predicate/function is stably computable by a CRD/CRC if and only if it is semilinear.
Definition 2.4.
$T\subseteq\mathbb{N}^{d}$ is a threshold set is if there are constants $c,w_{1},\dots,w_{d}\in\mathbb{Z}$ such that $T=\{\mathbf{x}\in\mathbb{N}^{d}\mid w_{1}\mathbf{x}(1)+\dots+w_{d}\mathbf{x}(d)\leq c\}.$ $M\subseteq\mathbb{N}^{d}$ is a mod set if there are constants $c,m,w_{1},\dots,w_{d}\in\mathbb{N}$ such that $M=\{\mathbf{x}\in\mathbb{N}^{d}\mid w_{1}\mathbf{x}(1)+\dots+w_{d}\mathbf{x}(d)\equiv c\mod m\}.$
The following well-known characterization of semilinear sets is useful.
Theorem 2.5 ([Ginsburg1966Semigroups]).
A set is semilinear if and only if it is a Boolean combination (union, intersection, complement) of threshold and mod sets.
3 Execution bounded chemical reaction networks
In this section, we define execution bounded CRNs and state some alternate characterizations of the definition.
Definition 3.1.
A CRN $\mathcal{C}$ is execution bounded from configuration $\mathbf{x}$ if all executions $\mathcal{E}=(\mathbf{x},\ldots)$ starting at $\mathbf{x}$ are finite. A CRD or CRC $\mathcal{C}$ is execution bounded if it is execution bounded from every valid initial configuration. $\mathcal{C}$ is entirely execution bounded if it is execution bounded from every configuration.
This is a distinct concept from the notion of “bounded” CRNs studied by Rackoff [Rackoff1978CoveringBoundedness] (studied under the equivalent formalism of vector addition systems). That paper defines a CRN to be bounded from a configuration $\mathbf{x}$ if $|\mathsf{reach}(\mathbf{x})|$ is finite (and shows that the decision problem of determining whether this is true is $\mathsf{EXPSPACE}$-complete.) We use the term execution bounded to avoid confusion with this concept.
We first observe an equivalent characterization of execution bounded that will be useful in the negative results of Section˜6.
Definition 3.2.
A execution $\mathcal{E}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots)$ is self-covering if for some $i<j$, $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$. It is strictly self-covering if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$. We also refer to these as (strict) self-covering paths. [^1] Rackoff [Rackoff1978CoveringBoundedness] uses the term “self-covering” to mean what we call strictly self-covering here, and points out that Karp and Miller [karp1969parallel] showed that $|\mathsf{reach}(\mathbf{x})|$ is infinite if and only if there is a strictly self-covering path from $\mathbf{x}$. The distinction between these concepts is illustrated by the CRN $A\mathop{\rightleftharpoons}\limits B$. From any configuration $\mathbf{x}$, $\mathsf{reach}(\mathbf{x})$ is finite ($|\mathsf{reach}(\mathbf{x})|=\mathbf{x}(A)+\mathbf{x}(B)+1$), and there is no strict self-covering path. However, from (say) $\{A\}$, there is a (nonstrict) self-covering path $\{A\}\Rightarrow\{B\}\Rightarrow\{A\}$, and by repeating, this CRN has an infinite cycling execution within its finite configuration space $\mathsf{reach}(\{A\})=\{\{A\},\{B\}\}$.
Lemma 3.3.
lem:execution-bdd-iff-no-self-covering-path A CRN is execution bounded from $\mathbf{x}$ if and only if there is no self-covering path from $\mathbf{x}$.
Proof 3.4.
For (the contrapositive of) the forward direction, assume there is a self-covering path from $\mathbf{x}$, which reaches to $\mathbf{x}_{i}$ and later to $\mathbf{x}_{j}\geqq\mathbf{x}_{i}$. By additivity the reactions leading from $\mathbf{x}_{i}$ to $\mathbf{x}_{j}$ can be repeated indefinitely (in a cycle if $\mathbf{x}_{i}=\mathbf{x}_{j}$, and increasing some molecular counts unboundedly if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$), so $\mathcal{C}$ is not execution bounded from $\mathbf{x}$.
For the reverse direction, assume $\mathcal{C}$ is not execution bounded from $\mathbf{x}$. Then there is an infinite execution $\mathcal{E}=(\mathbf{x}=\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\dots).$ By Dickson’s Lemma there are $i<j$ such that $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$, i.e., $\mathcal{E}$ is self-covering.
4 Execution bounded CRDs stably decide all semilinear sets
In this section, we will show that execution bounded CRDs have the same computational power as unrestricted CRDs. The following is the main result of this section.
Theorem 4.1.
Exactly the semilinear sets are stably decidable by execution bounded CRDs. Furthermore, each can be stably decided with expected stabilization time $\Theta(n\log n)$.
Since semilinear sets are Boolean combinations of mod and threshold predicates, we prove this theorem by showing that execution bounded CRDs can decide mod and threshold sets individually as well as any Boolean combination in the following lemmas. To ensure execution boundedness in the last step, we require the following property.
Definition 4.2.
Let $\mathcal{D}$ be a CRD with voting species $\Upsilon$. We say $\mathcal{D}$ is single-voting if for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Sigma}$ and any $\mathbf{c}\in\mathbb{N}^{\Lambda}$ s.t. $\mathbf{i}\Rightarrow\mathbf{c}$, $\sum_{V\in\Upsilon}\mathbf{c}(V)=1$, i.e., exactly one voter is present in every reachable configuration.
Lemma 4.3.
lem:mod_sets_decidable Every mod set $M=\big\{\left(x_{1},\ldots,x_{d}\right)\mid\sum_{i=1}^{d}w_{i}x_{i}\equiv c\bmod m\big\}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
We design a CRD $\mathcal{D}$ with exactly one leader present at all times, cycling through $m$ “states” while consuming the input and accepting on state $c$. Let $\Sigma=\{X_{1},\dots,X_{d}\}$ be the set of input species and start with only one $L_{0}$ leader, i.e. set the initial context $\mathbf{s}(L_{0})=1$ and $\mathbf{s}(S)=0$ for all other species. For each $i\in\{1,\dots,d\},j\in\{0,\dots,m-1\}$ add the following reaction: $X_{i}+L_{j}\rightarrow L_{j+w_{i}\bmod m}.$ Let only $L_{c}$ vote yes and all other species no, i.e. $\Upsilon=\{L_{c}\}$. For any valid initial configuration, $\mathcal{D}$ reaches a stable configuration which votes yes if and only if the input is in the mod set, and no otherwise.
Proof 4.4.
$\mathcal{D}$ terminates with the correct output value: At any point in time, there is a single leader $L_{j}$ present (the initial configuration contains a single leader and each reaction produces and consumes one). Every reaction satisfies the following invariant (for the leader’s subscript $j$): $j\equiv\sum_{i=1}^{d}w_{i}x_{i}^{\prime}\bmod m$ where $x_{i}^{\prime}$ is the updated count of species $X_{i}$ in the current configuration. By design of $\mathcal{D}$, there will be a reaction applicable as long as there are copies of $X_{i}$ (a leader with any subscript can react with any $X_{i}$). After applying this reaction as often as possible, we have reached a stable configuration with $L_{\sum_{i=1}^{d}w_{i}x_{i}\bmod m}$ as the only species present.
$\mathcal{D}$ is execution bounded: Every reaction reduces the count of chemicals by one. Every possible execution contains exactly $\|\mathbf{i}\|$ configurations, where $\|\mathbf{i}\|$ is the number of all molecules in the starting configuration.
$\mathcal{D}$ is single-voting: Initially, $L_{0}$ is present and the only voter. Every valid input contains no voter and every reaction results in no change to the count of copies of $L_{i}$.
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: We start with #L=1,#X=n\#L=1,\#X=n in volume $n$. $n$ reactions must occur before $\mathcal{D}$ terminates. For the first reaction, we have a rate of $\lambda=\frac{n\cdot 1}{n}$, for the last (with only the leader and one $X$ present), our rate will be $\lambda=\frac{1\cdot 1}{n}$. Thus, the expected time for all $n$ reactions to complete is
$\sum_{i=1}^{n}\frac{n}{i}=n\sum_{i=1}^{n}\frac{1}{i}=\Theta(n\log n).$
Lemma 4.5.
lem:threshold_sets_decidable Every threshold set $T=\big\{(x_{1},\ldots,x_{d})\mid\sum_{i=1}^{d}w_{i}x_{i}\geq t\big\}$ is stably decidable by an execution bounded, single-voting CRD with expected stabilization time $\Theta(n\log n)$.
We design a CRD $\mathcal{D}$ which multiplies the input molecules according to their weight and consumes positive and negative units alternatingly using a single leader. Once no more reaction is applicable, the leader’s state will indicate whether or not there are positive units left and the threshold is met. Let $\Sigma=\{X_{1},\dots,X_{d}\}$ be the set of input species and $\Upsilon=\{L_{Y}\}$ the yes voter. We first add reactions to multiply the input species by their respective weights. For all $i\in\{1,\dots,d\}$, add the reaction:
$X_{i}\rightarrow\begin{cases}w_{i}P&\text{if }w_{i}>0\\ -w_{i}N&\text{if }w_{i}<0\\ \emptyset&\text{otherwise}\end{cases}$ (5)
$P$ and $N$ represent “positive” and “negative” units respectively. Now add reactions to consume $P$ and $N$ alternatingly using a leader until we run out of one species:
$L_{Y}+N$ $\rightarrow L_{N}$ (6)
$L_{N}+P$ $\rightarrow L_{Y}$ (7)
Finally, initialize the CRD with one $L_{Y}$ and the threshold number $t$ copies of $P$ (or tN-tN if $t$ is negative), i.e. $\mathbf{s}(L_{Y})=1$, $\mathbf{s}(P)=t$ if $t>0$, or $\mathbf{s}(N)=-t$ if $t<0$, and $\mathbf{s}(S)=0$ for all other species. For any valid initial configuration, $\mathcal{D}$ reaches a stable configuration which votes yes if and only if the weighted sum of inputs is above the threshold, and no otherwise.
Proof 4.6.
$\mathcal{D}$ is single-voting since it starts with a single leader and no reaction changes the count of $L_{B}$ molecules.
$\mathcal{D}$ stabilizes in $\Theta(n\log n)$ time: First, all input species will be converted to $w_{i}$ instances of $P$ or $N$. We run these reactions until no $X_{i}$. As they are independent of molecules other than the reactant, these reactions have a rate of $\lambda=i$, so the expected time until the next reaction is $\frac{1}{i}$. The total time for reactions (5) to complete is therefore $\sum_{i=1}^{n}\frac{1}{i}=\mathit{\Theta}(\log n)$. The time for reactions (6) and (7) on the other hand is asymptotically dominated by the last reaction, where #L=1\#L=1 and #B=1\#B=1, where $B\in\{P,N\}$, so $\lambda=\frac{1\cdot 1}{n}$. Let $n_{P},n_{N}$ be the counts of $P,N$ and assume without loss of generality $n_{P}\geq n_{N}$. We get:
$\sum_{i=0}^{n_{N}-1}\frac{n}{(n_{P}-i)}+\frac{n}{(n_{N}-i)}\leq\sum_{i=0}^{n_{N}-1}\frac{2n}{(n_{N}-i)}=2n\sum_{i=1}^{n_{N}}\frac{1}{i}=\Theta(n\log n).$
Lemma 4.7.
If sets $X_{1},X_{2}\subseteq\mathbb{N}^{d}$ are stably decided by some execution bounded, single-voting CRD, then so are $X_{1}\cup X_{2},X_{1}\cap X_{2}$, and $\overline{X_{1}}$ with expected stabilization time $O(n\log n)$.
Proof 4.8.
To stably decide $\overline{X_{1}}$, swap the yes and no voters.
For $\cup$ and $\cap$, consider a construction where we decide both sets separately and record both of their votes in a new voter species. For this, we allow the set of all voters to be a strict subset of all species. We first add reactions to duplicate our input with reactions of the form
$X_{i}\rightarrow X_{i,1}+X_{i,2}$ (8)
by two separate CRDs. Subsequently, we add reactions to record the separate votes in one of four new voter species: $V_{NN},V_{NY},V_{YN},V_{YY}$. The first and second CRN determine the first and second subscript respectively. For $b\in\{Y,N\}$ and if $S_{b},T_{b}$ are voters of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, add the reactions:
$S_{b}+V_{\overline{b}?}\rightarrow S_{b}+V_{b?}$ (9)
$T_{b}+V_{?\overline{b}}\rightarrow T_{b}+V_{?b}$ (10)
Above, the ? subscript is shorthand for “any bit”; e.g. if $N_{1}$ is the no voter of the first CRD, we would add two reactions $N_{1}+L_{YN}\rightarrow N_{1}+L_{NN}$ and $N_{1}+L_{YY}\rightarrow N_{1}+L_{NY}$. We let the yes voters be: $\Upsilon=\{V_{NY},V_{YN},V_{YY}\}$ to stably decide $X_{1}\cup X_{2}$ or $\Upsilon=\{V_{YY}\}$ to stably decide $X_{1}\cap X_{2}$.
Reaction (8) will complete in $O(\log n)$ time and is clearly execution bounded since the input $X_{i}$ is finite and not produced in any reaction. Consequently, two separate CRNs run in $\mathit{\Theta}(n\log n)$ time as shown in LABEL:lem:mod_sets_decidable and LABEL:lem:threshold_sets_decidable. After stabilization of the parallel CRNs, we expect reaction (9) and (10) to happen exactly once. Each molecule involved is a leader and has count $1$ in volume $n$. This leads to a rate of $\lambda=\frac{1\cdot 1}{n}$, so the expected time for one reaction to happen is $O(n)$. It is important to note that reactions (9) and (10) do not result in unbounded executions due to the unanimous vote in parallel CRDs. In both mod sets and threshold sets, the leader changes its vote a maximum of $|\mathbf{i}|$ times, with only ever one leader present at any time. Again, we start with only one $V_{bb}$ voter present initially and no reaction changes the count of voters, making our construction single-voting.
Since semilinear predicates are exactly Boolean combinations of threshold and mod predicates, LABEL:lem:mod_sets_decidable, LABEL:lem:threshold_sets_decidable and 4.7 imply Theorem˜4.1.
We can also prove the same result for all-voting CRDs. Note, however, that such CRDs cannot be “composed” using the constructions of Lemmas˜4.7 and 5.6, which crucially relied on the assumption that the CRDs being used as “subroutines” are single-voting.
Theorem 0.
lem:boolean-closure-crd-all-voting Every semilinear set is stably decidable by an execution bounded, all-voting CRD, with expected stabilization time $O(n\log n).$
Proof 4.10.
[…]
$V^{b}+S^{\overline{b}}\mathop{\rightarrow}\limits V^{b}+S^{b}$ (11)
The original reactions of the CRD must also be replaced with “functionally identical” reactions for the new versions of species. For example, the reaction $A+B\mathop{\rightarrow}\limits C+D$ becomes
$A^{0}+B^{0}$ $\mathop{\rightarrow}\limits C^{0}+D^{0}$
$A^{0}+B^{1}$ $\mathop{\rightarrow}\limits C^{0}+D^{0}$
$A^{1}+B^{0}$ $\mathop{\rightarrow}\limits C^{0}+D^{0}$
$A^{1}+B^{1}$ $\mathop{\rightarrow}\limits C^{1}+D^{1}$
In the middle two cases we can pick the superscripts of the products arbitrarily, whereas in the first and last case, we must choose the product votes to match those of the reactants to ensure stable states remain stable.
A vote change of the $V^{b}$ leader leads to the propagation of the vote to at most $n$ molecules once using reaction (11). This reaction dominates the runtime, as a single molecule is required to interact with each other molecule. We cannot speed this process up using an epidemic style process as conflicting votes would make the CRN execution unbounded. The original CRD takes time $O(n\log n)$ to converge on a correct output for the single voter $V^{b}$. At that point, a standard coupon collector argument shows that the voter $V^{b}$ takes expected time $O(n\log n)$ to correct the votes of all other species via reaction (11).
5 Execution bounded CRCs stably compute all semilinear functions
In this section we shift focus from computing Boolean-valued predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ to integer-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}$, showing that execution bounded CRCs can stably compute the same class of functions (semilinear) as unrestricted CRCs.
Similar to [Chen2012DeterministicFunction, Doty2013Leaderless], we compute semilinear functions by decomposing them into “affine pieces”, which we will show can be computed by execution bounded CRNs and combined by using semilinear predicates to decide which linear function to apply for a given input. [^2] While this proof generalizes to multivariate output functions as in [Chen2012DeterministicFunction, Doty2013Leaderless], to simplify notation we focus on single output functions. Multi-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}^{l}$ can be equivalently thought of as $l$ separate single output functions $f_{i}:\mathbb{N}^{d}\to\mathbb{N}$, which can be computed in parallel by independent CRCs.
We say a partial function $f:\mathbb{N}^{k}\dashrightarrow\mathbb{N}$ is affine if there exist a vectors $\mathbf{a}\in\mathbb{Q}^{k}$, $\mathbf{c}\in\mathbb{N}^{k}$ with $\mathbf{x}-\mathbf{c}\geq\mathbf{0}$ and nonnegative integer $b\in\mathbb{N}$ such that $f(\mathbf{x})=\mathbf{a}^{\top}(\mathbf{x}-\mathbf{c})+b.$ For a partial function $f$ we write $\operatorname{dom}f$ for the domain of $f$, the set of inputs for which $f$ is defined. This definition of affine function may appear contrived, but the main utility of the definition is that it satisfies LABEL:lem:partial_affine_function_decomposition_disjoint_domains. For convenience, we can ensure to only work with integer valued molecule counts by multiplying by $\frac{1}{d}$ after the dot product, where $d$ may be taken to be the least common multiple of the denominators of the rational coefficients in the original definition such that $n_{i}=d\cdot\mathbf{a}(i)$:
$f(\mathbf{x})=b+\sum_{i=1}^{k}a_{i}(x_{i}-c_{i})\iff f(\mathbf{x})=b+\frac{1}{d}\sum_{i=1}^{k}n_{i}(x_{i}-c_{i}).$
We say that a partial function $\hat{f}:\mathbb{N}^{k}\rightarrow\mathbb{N}^{2}$ is a diff-representation of $f$ if $\operatorname{dom}f=\operatorname{dom}\hat{f}$ and, for all $\mathbf{x}\in\operatorname{dom}f$, if $\left(y_{P},y_{C}\right)=\hat{f}(\mathbf{x})$, then $f(\mathbf{x})=y_{P}-y_{C}$, and $y_{P}=O(f(\mathbf{x}))$. In other words, $\hat{f}$ represents $f$ as the difference of its two outputs $y_{P}$ and $y_{C}$, with the larger output $y_{P}$ possibly being larger than the original function’s output, but at most a multiplicative constant larger [Doty2013Leaderless].
Lemma 5.1.
Let $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ be an affine partial function. Then there is a diff-representation $\hat{f}:\mathbb{N}^{k}\longrightarrow\mathbb{N}^{2}$ of $f$ and an execution bounded CRC that monotonically stably computes $\hat{f}$ in expected stabilization time $O(n)$.
Proof 5.2.
Define a CRC $C$ with input species $\Sigma=\{X_{1},\ldots,X_{k}\}$ and output species $\Gamma=\{Y^{P},Y^{C}\}$. We need to ensure that after stabilizing, $y=\#Y^{P}-\#Y^{C}$
To account for the $b$ offset, start with $b$ copies of $Y^{P}$.
For the $c_{i}$ offset, we must reduce the number of $X_{i}$ by $c_{i}$. Since the result will be used in the next reaction, we want to produce a new species $X_{i}^{\prime}$ and require $X_{i}^{\prime}$ to not be consumed during the computation. We achieve this by adding reactions that let $X_{i}$ consume itself $c_{i}$ times (keeping track with a subscript) and converting $X_{i}$ to $X_{i}^{\prime}$ once $c_{i}$ has been reached. For the sake of notation below, assume input species $X_{i}$ is actually named $X_{i,1}$. For each $i\in\{1,\ldots,k\}$ and $m,p\in\left\{1,\ldots,c_{i}\right\}$, if $m+p\leq c_{i}$, add the reaction
$X_{i,m}+X_{i,p}\rightarrow X_{i,m+p}$ (12)
If $m+p>c_{i}$, add the reaction
$X_{i,m}+X_{i,p}\rightarrow X_{i,c_{i}}+\left(m+p-c_{i}\right)X_{i}^{\prime}$ (13)
Runtime: In volume $n$, the rate of reactions (12) and (13) would be $\lambda\approx\frac{(x_{i})^{2}}{n}$ ($x_{i}$ molecules have the chance to react with any of the $x_{i}-1$ others), so the expected time for the next reaction is $\frac{n}{(x_{i})^{2}}$. The expected time for the whole process is $\sum_{i=1}^{x_{i}}\frac{n}{i^{2}}=n\sum_{i=1}^{x_{i}}\frac{1}{i^{2}}=O(n)$. Further, the reactions are execution bounded since both strictly decrease the number of their reactants and exactly $x_{i}-1$ reactions will happen.
To account for the $n_{i}/d$ coefficient, we multiply by $n_{i}$, then divide by $d$ using similar reactions as for the subtraction. To multiply by $n_{i}$, add the following reaction for each $i\in\{1,\ldots,k\}$:
$X_{i}^{\prime}\rightarrow\begin{cases}n_{i}D_{1}^{P},&\text{ if }n_{i}>0\\ \left(-n_{i}\right)D_{1}^{C},&\text{ if }n_{i}<0\end{cases}$ (14)
For each $m,p\in\left\{1,\ldots,d-1\right\}$, if $m+p\leq d-1$, add the reactions
$D_{m}^{P}+D_{p}^{P}\rightarrow D_{m+p}^{P}$ (15)
$D_{m}^{C}+D_{p}^{C}\rightarrow D_{m+p}^{C}$ (16)
If $m+p>c_{i}$, add the reactions
$D_{m}^{P}+D_{p}^{P}\rightarrow D_{m+p-d}^{B}+Y^{P}$ (17)
$D_{m}^{C}+D_{p}^{C}\rightarrow D_{m+p-d}^{B}+Y^{C}$ (18)
Reactions (14) complete in expected time $O(\log n)$, while (17) and (18) complete in $O(n)$ by a similar analysis as for the first two reactions. As for execution boundedness, (14) is only applicable once for every $X_{i}^{\prime}$; all other reactions start with a number of reactants which are a constant factor of $X_{i}^{\prime}$ and decrease the count of their reactants by one in each reaction.
We require the following result due to Chen, Doty, Soloveichik [Chen2012DeterministicFunction], guaranteeing that any semilinear function can be built from affine partial functions.
Lemma 5.3 ([Chen2012DeterministicFunction]).
Let $f:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a semilinear function. Then there is a finite set $\big\{f_{1}:\mathbb{N}^{d}\rightarrow\mathbb{N},\ldots,f_{m}:\mathbb{N}^{d}\rightarrow\mathbb{N}\big\}$ of affine partial functions, where each $\operatorname{dom}f_{i}$ is a linear set, such that, for each $\mathbf{x}\in\mathbb{N}^{d}$, if $f_{i}(\mathbf{x})$ is defined, then $f(\mathbf{x})=f_{i}(\mathbf{x})$, and $\bigcup_{i=1}^{m}\operatorname{dom}f_{i}=\mathbb{N}^{d}$.
We strengthen Lemma˜5.3 to show we may assume each $\mathrm{dom}\ f_{i}$ is disjoint from the others. This is needed not only to prove Theorem˜5.6, but to correct the proof of Lemma 4.4 in [Chen2012DeterministicFunction], which implicitly assumed the domains are disjoint.
Lemma 5.4.
lem:partial_affine_function_decomposition_disjoint_domains Let $f:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a semilinear function. Then there is a finite set $\big\{f_{1}:\mathbb{N}^{d}\rightarrow\mathbb{N},\ldots,f_{m}:\mathbb{N}^{d}\rightarrow\mathbb{N}\big\}$ of affine partial functions, where each $\operatorname{dom}f_{i}$ is a linear set, and $\operatorname{dom}f_{i}\cap\operatorname{dom}f_{j}=\emptyset$ for all $i\neq j$, such that, for each $\mathbf{x}\in\mathbb{N}^{d}$, if $f_{i}(\mathbf{x})$ is defined, then $f(\mathbf{x})=f_{i}(\mathbf{x})$, and $\bigcup_{i=1}^{m}\operatorname{dom}f_{i}=\mathbb{N}^{d}$.
Proof 5.5.
By [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2], every semilinear set is a finite union of disjoint fundamental linear sets. The author defines a linear set $L=\big\{\mathbf{b}+n_{1}\mathbf{u}_{1}+\ldots+n_{p}\mathbf{u}_{p}\mid n_{1},\ldots,n_{p}\in\mathbb{N}\big\}$ as fundamental, if $\mathbf{u}_{1},\dots\mathbf{u}_{p}\in\mathbb{N}^{k}$ span a $p$-dimensional vector space in $\mathbb{R}^{k}$, i.e. all vectors are linearly independent in $\mathbb{R}^{k}$. [^3] This distinction is significant because not all integer-valued linear sets can be represented using solely linearly independent vectors. An illustrative example is $\mathbf{b}=\mathbf{0},\mathbf{u}_{1}=(1,1,1),\mathbf{u}_{2}=(2,0,1),\mathbf{u}_{3}=(0,2,1)$, as discussed in [Chen2012DeterministicFunction]. The vectors $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}$ are not linearly independent in $\mathbb{R}^{3}$, yet this set cannot be expressed with less than three basis vectors. The proof of Lemma˜5.3 in [Chen2012DeterministicFunction] shows that each linear set $L_{i}$ comprising the semilinear graph of $f$ corresponds to one partial affine function $f_{i}$. The fact that [Ito1969SemilinearSetsFiniteUnionDisjointLinearSets, Theorem 2] lets us assume each $L_{i}$ is disjoint from the others immediately implies that each $\mathrm{dom}\ f_{i}$ is disjoint from the others.
The next theorem shows that semilinear functions can be computed by execution bounded CRCs in expected time $O(n\log n)$.
Theorem 5.6.
Let $f:\mathbb{N}^{d}\to\mathbb{N}$ be a semilinear function. Then there is an execution bounded CRC that stably computes $f$ with expected stabilization time $O(n\log n)$.
Proof 5.7.
We employ the same construction of [Chen2012DeterministicFunction] with minor alterations. A CRC with input species $\Sigma=\{X_{1},\ldots,X_{d}\}$ and output species $\Gamma=\{Y\}$. By LABEL:lem:partial_affine_function_decomposition_disjoint_domains, we decompose our semilinear function into partial affine functions (with linear, disjoint domains), which can be computed in parallel by Lemma˜5.1. Further, we decide which function to use by computing the predicate $\phi_{i}=$$\mathrm{x}\in\operatorname{dom}f_{i}$” (Theorem˜4.1). We interpret each $\widehat{Y}_{i}^{P}$ and $\widehat{Y}_{i}^{C}$ as an “inactive” version of “active” output species $Y_{i}^{P}$ and $Y_{i}^{C}$. Let $L_{i}^{Y},L_{i}^{N}$ be the yes and no voters respectively voting whether $\mathbf{x}$ lies in the domain of $i$-th partial function. Now, we convert the function result of the applicable partial affine function to the global output by adding the following reactions for each $i\in\{1,\ldots,m\}$.
$L_{i}^{Y}+\widehat{Y}_{i}^{P}$ $\rightarrow L_{i}^{Y}+Y_{i}^{P}+Y$ (19)
$L_{i}^{N}+Y_{i}^{P}$ $\rightarrow L_{i}^{N}+M_{i}$ (20)
$M_{i}+Y$ $\rightarrow\widehat{Y}_{i}^{P}$ (21)
Reaction (19) produces an output copy of species $Y$ and (20) and (21) reverse the first reaction using only bimolecular reactions. Both are catalyzed by the vote of the $i$-th predicate result. Also add reactions
$L_{i}^{Y}+\widehat{Y}_{i}^{C}\rightarrow L_{i}^{Y}+Y_{i}^{C}$ (22)
$L_{i}^{N}+Y_{i}^{C}\rightarrow L_{i}^{N}+\widehat{Y}_{i}^{C}$ (23)
and
$Y_{i}^{P}+Y_{i}^{C}$ $\rightarrow K$ (24)
$K+Y$ $\rightarrow\varnothing$ (25)
Reactions (22) and (23) activate and deactivate the “negative” output values and reactions (24) and (25) allow two active partial outputs to cancel out and consume the excess $Y$ in the process. When the input is in the domain of function $i$, exactly one copy of $L_{i}^{Y}$ will be present, otherwise one copy of $L_{i}^{N}$. Since we know that the predicate computation is execution bounded and produces at most one voter, the catalytic reaction will also happen at most as often as the leader changes its vote. Therefore, it is also execution bounded.
The underlying CRNs computing the predicates and functions have expected stabilization time $O(n\log n).$ Once they have stabilized, the slowest reactions described above are those where a leader ($L_{i}^{Y}$ or $L_{i}^{N}$) must convert all outputs, which also takes expected time $O(n\log n)$ by a coupon collector argument.
6 Limitations of execution bounded CRNs
The main positive results of the paper (Theorems˜4.1 and 5.6) rely on the assumption that valid initial configurations have a single leader (in particular, they are execution bounded only from configurations with a single leader, but not from arbitrary configurations). LABEL:lem:boolean-closure-crd-all-voting shows that we may assume the CRD deciding a semilinear set is all-voting. However, for the “constructive” results LABEL:{lem:boolean-closure-crd} and 5.6, which compose the output of a CRD $\mathcal{D}$ with downstream computation, using $\mathcal{D}$ as a “subroutine” to stably compute a more complex set/function, the constructions crucially use the assumption that $\mathcal{D}$ is single-voting (i.e., only the leader of $\mathcal{D}$ votes) to argue the resulting composed CRN is execution bounded. In this section we show these assumptions are necessary, proving that execution bounded CRNs without those constraints are severely limited in their computational abilities.
We show that entirely execution bounded CRNs (from every configuration) can be characterized by a simpler property of having a “linear potential function” that essentially measures how close the CRN is to reaching a terminal configuration. This result was essentially proven independently by Czerner, Guttenberg, Helfrich, and Esparza [czerner2024fast] in the model of population protocols. We give a self-contained proof since we technically need the more general CRN model, although our proof appears to follow their same technique of using a Farkas-like lemma. We use this characterization to prove that entirely execution bounded CRNs can stably decide only limited semilinear predicates (eventually constant, Definition˜6.9), assuming all species vote, and that molecular counts cannot decrease to $O(1)$ in stable configurations (see Definition˜6.6).
6.1 Linear potential functions
We define a linear potential function of a CRN to be a nonnegative linear function of configurations that each reaction strictly decreases.
Definition 6.1.
A linear potential function $\Phi:\mathbb{R}_{\geq 0}^{\Lambda}\to\mathbb{R}_{\geq 0}$ for a CRN is a nonnegative linear function, such that for each reaction $(\mathbf{r},\mathbf{p})$, $\Phi(\mathbf{p})-\Phi(\mathbf{r})<0.$
Note that for a configuration $\mathbf{x}$, since $\Phi(\mathbf{x})=\sum_{S\in\Lambda}v_{S}\mathbf{x}(S)\geq 0$, it must be nondecreasing in each species, i.e., all coefficients $v_{S}$ must be nonnegative (though some are permitted to be 0). Intuitively, we can think of $\Phi$ as assigning a nonnegative “mass” to each species (the mass of $S$ is $v_{S}$), such that each reaction removes a positive amount of mass from the system. Note also that since $\Phi$ is linear, the above is equivalent to requiring that $\Phi(\mathbf{p}-\mathbf{r})<0$, if we extend $\Phi$ to a linear function $\Phi:\mathbb{R}^{\Lambda}\to\mathbb{R}$ on vectors with negative elements.
A CRN may or may not have a linear potential function. Although it is not straightforward to “syntactically check” a CRN to see if has a linear potential function, it is efficiently decidable: a CRN has a linear potential function if and only if the following system of linear inequalities has a solution (which can be solved in polynomial time using linear programming techniques; the variables to solve for are the $v_{S}$ for each $S\in\Lambda$), where the $i$’th reaction has reactants $\mathbf{r}_{i}$ and products $\mathbf{p}_{i}$, and species $S\in\Lambda$ has mass $v_{S}\geq 0$:
$(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]v_{S}<0$
For example, for the reactions $A+A\mathop{\rightarrow}\limits B+C$ and $B+B\mathop{\rightarrow}\limits A$, for each reaction to strictly decrease the potential function $\Phi(\mathbf{x})=v_{A}\mathbf{x}(A)+v_{B}\mathbf{x}(B)+v_{C}\mathbf{x}(C)$, $\Phi$ must satisfy $2v_{A}>v_{B}+v_{C}$ and $2v_{B}>v_{A}$. In this case, $v_{A}=1,v_{B}=1,v_{C}=0$ works.
Remark 6.2.
A system of linear inequalities with rational coefficients has a real solution if and only if it has a rational solution. For any homogeneous system (where all inequalities are comparing to 0), any positive scalar multiple of a solution is also a solution. By clearing denominators, a system has a rational solution if and only if it has an integer solution. Thus, one can equivalently define a linear potential function to be a function $\Phi(\mathbf{x})=\sum_{S\in\Lambda}v_{S}\mathbf{x}(S)$ such that each $v_{S}\in\mathbb{N}$, i.e., we may assume $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$. In particular, since $\Phi$ is decreased by each reaction, it is decreased by at least 1.
The following is a variant of Farkas’ Lemma [farkas1902theorie], one of several similar “Theorems of the Alternative” stating that exactly one of two different linear systems has a solution. (See [mangasarian1994nonlinear, Chapter 2.4] for a list of such theorems.) A proof can be found in [gale1960theory, Theorem 2.10].
Theorem 6.3.
Let $\mathbf{M}$ be a real matrix. Exactly one of the following statements is true.
  1. 1.
    There is a vector $\mathbf{u}\geqq\mathbf{0}$ such that $\mathbf{M}\mathbf{u}<\mathbf{0}$.
  2. 2.
    There is a vector $\mathbf{v}\geq\mathbf{0}$ such that $\mathbf{v}\mathbf{M}\geqq\mathbf{0}$.
We require the following discrete variant of Theorem˜6.3. The geometric intuition of this version is illustrated in Figure˜1.
Corollary 6.4.
cor:linear-alternative-integer Let $\mathbf{M}$ be a rational matrix. Exactly one of the following statements is true.
  1. 1.
    There is an integer vector $\mathbf{u}\geq\mathbf{0}$ such that $\mathbf{M}\mathbf{u}\geqq\mathbf{0}$.
  2. 2.
    There is an integer vector $\mathbf{v}\geqq\mathbf{0}$ such that $\mathbf{v}\mathbf{M}<\mathbf{0}$.
Proof 6.5.
For convenience when we use LABEL:cor:linear-alternative-integer in proving LABEL:thm:potential-function-iff-entirely-execution-bounded, we swapped the roles of $\mathbf{u}$ and $\mathbf{v}$ in left- vs. right-multiplication with $\mathbf{M}$; the real-valued version of the statement of LABEL:cor:linear-alternative-integer is equivalent to Theorem˜6.3 by taking the transpose of $\mathbf{M}$.
To see that we may assume the vectors are integer-valued if $\mathbf{M}$ is rational-valued, recall that a system of linear equalities/inequalities with rational coefficients has a solution if and only if it has a rational solution. Since the system is homogeneous (the matrix-vector product is compared to the zero vector $\mathbf{0}$), any multiple of a solution is also a solution. By clearing denominators, it has a rational solution if and only if it has an integer solution.
Figure 1: Geometric intuition of LABEL:cor:linear-alternative-integer. A matrix $\mathbf{M}$ has column vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. The cone of $\mathbf{M}$ is the nonnegative span of these vectors, shown as a faded gray region. Exactly one of two scenarios occurs: a) The cone of $\mathbf{M}$ intersects the first quadrant (nonnegative orthant in higher dimensions) away from the origin, i.e., some semipositive point ($\mathbf{M}\mathbf{u}\geq\mathbf{0}$ above) is a nonnegative linear combination of vectors $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}$. b) The cone of $\mathbf{M}$ does not intersect the first quadrant except at the origin. In this case we can draw a dashed line (hyperplane in higher dimensions) separating the cone of $\mathbf{M}$ from the first quadrant. The orthogonal vector $\mathbf{v}\geq\mathbf{0}$ to this line lies in the first quadrant, but $\mathbf{v}\mathbf{M}<0$ means each vector $\mathbf{x}_{i}$ has negative dot product $\mathbf{v}\cdot\mathbf{x}_{i}<0$, i.e., all $\mathbf{x}_{i}$ vectors lie on the other side of the line.
6.2 Impossibility of stably deciding majority and parity
In this section, we prove Theorem˜6.7, which is a special case of our main negative result, LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant. We give a self-contained proof of Theorem˜6.7 because it is simpler and serves as an intuitive warmup to some of the key ideas used in proving LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant, without the complexities of dealing with arbitrary semilinear sets.
Theorem˜6.7 shows a limitation on the computational power of entirely execution bounded, all-voting CRNs, but it requires an additional constraint on the CRN for the result to hold (and we later give counterexamples showing that this extra hypothesis is provably necessary), described in the following definition.
Definition 6.6.
Let $\mathcal{D}$ be a CRD. The output size of $\mathcal{D}$ is the function $s:\mathbb{N}\to\mathbb{N}$ defined $s(n)=\min_{\mathbf{x},\mathbf{y}}\{\|\mathbf{y}\|\mid\mathbf{x}\Rightarrow\mathbf{y},\|\mathbf{x}\|=n,\mathbf{x}\text{ is a valid initial configuration},\mathbf{y}\text{ is stable}\}$, the size of the smallest stable configuration reachable from any valid initial configuration of size $n$. A CRD is non-collapsing if $\lim_{n\to\infty}s(n)=\infty$.
Put another way, $\mathcal{D}$ is collapsing if there is a constant $c$ such that, from infinitely many initial configurations $\mathbf{x}$, $\mathcal{D}$ can reach a stable configuration of size at most $c$. All population protocols are non-collapsing, since every reaction preserves the configuration size.
Theorem 6.7.
No non-collapsing, all-voting, entirely execution bounded CRD can stably decide the majority predicate $X_{1}\geq X_{2}?]$ or the parity predicate $X\equiv 1\mod 2?]$.
Proof 6.8.
Let $\mathcal{D}=(\Lambda,R,\Sigma,\Upsilon_{\mathrm{Y}},\Upsilon_{\mathrm{N}},\mathbf{s})$ be a CRD obeying the stated conditions, and suppose for the sake of contradiction that $\mathcal{D}$ stably decides the majority predicate (so $\Sigma=\{X_{1},X_{2}\}$).
We consider the sequence of stable configurations $\mathbf{a}_{1},\mathbf{b}_{1},\mathbf{a}_{2},\mathbf{b}_{2},\dots$ defined as follows. Let $\mathbf{a}_{1}$ be a stable configuration reachable from initial configuration $\mathbf{s}+\{X_{1},X_{2}\}$; since the correct answer is yes, all species present in $\mathbf{a}_{1}$ vote yes. Now add a single copy of $X_{2}$. By additivity, the configuration $\mathbf{a}_{1}+\{X_{2}\}$ is reachable from $\mathbf{s}+\{X_{1},2X_{2}\}$, for which the correct answer in this case is no. Thus, since $\mathcal{D}$ stably decides majority, from $\mathbf{a}_{1}+\{X_{2}\}$, a stable “no” configuration is reachable; call this $\mathbf{b}_{1}$. Now add a single $X_{1}$. Since the correct answer is yes, from $\mathbf{b}_{1}+\{X_{1}\}$ a stable “yes” configuration is reachable, call it $\mathbf{a}_{2}$.
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{1},\mathbf{b}_{1},\mathbf{a}_{2},\mathbf{b}_{2},\dots$ where all species in $\mathbf{a}_{i}$ vote yes and all species in $\mathbf{b}_{i}$ vote no. Since $\mathcal{D}$ is non-collapsing, the size of the configurations $\mathbf{a}_{i}$ and $\mathbf{b}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically increasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.) Since all species vote, for some constant $\delta>0$, to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$, at least $\delta\|\mathbf{a}_{i}\|$ reactions must occur. This is because all species in $\mathbf{a}_{i}$ must be removed since they vote yes, and each reaction removes at most $O(1)$ molecules. (Concretely, let $\delta=1/\max_{(\mathbf{r},\mathbf{p})\in R}\|\mathbf{r}\|-\|\mathbf{p}\|$, i.e., 1 over the most net molecules consumed in any reaction.) Similarly, to get from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, at least $\delta\|\mathbf{b}_{i}\|$ reactions must occur.
Since $\mathcal{D}$ is entirely execution bounded, by LABEL:thm:potential-function-iff-entirely-execution-bounded, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$, where $\mathbf{v}\geq\mathbf{0}$. Adding a single $X_{2}$ to $\mathbf{a}_{i}$ increases $\Phi$ by the constant $\mathbf{v}(X_{2})$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $X_{2}$ and then going from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$). Similarly, adding a single $X_{1}$ to $\mathbf{b}_{i}$ and going from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, the resulting total change in $\Phi$ is unbounded and (for large enough $i$) negative.
$\Phi$ starts this process at the constant $\Phi(\mathbf{s}+\{X_{1},X_{2}\})$. Before $\|\mathbf{a}_{i}\|$ and $\|\mathbf{b}_{i}\|$ are large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$ and $\delta\|\mathbf{b}_{i}\|\geq\mathbf{v}(X_{1})+2$ (i.e., large enough that the net change in $\Phi$ is negative resulting from adding a single input and going to the next stable configuration), $\Phi$ could increase, if $\Phi(\{X_{1}\})$ (resp. $\Phi(\{X_{2}\})$) is larger than the net decrease in $\Phi$ due to following reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ (resp. from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i}$).
However, since $\mathcal{D}$ is non-collapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{s}+\{X_{1},X_{2}\})$), after which $\Phi$ strictly decreases after each round of this process. At some point in this process, $\mathcal{D}$ will not be able to reach all the way to the next $\mathbf{a}_{i}$ or $\mathbf{b}_{i}$ without $\Phi$ becoming negative, a contradiction.
The argument for parity is similar, but instead of alternating adding $X_{1}$ then $X_{2}$, in each round we always add one more $X$ to flip the correct answer.
Theorem˜6.7 is false without the non-collapsing hypothesis. The following collapsing, leaderless (but all-voting and entirely execution bounded) CRD stably decides majority: Species $X_{1},x_{1}$ vote yes, while $X_{2},x_{2}$ vote no:
$X_{1}+X_{2}$ $\mathop{\rightarrow}\limits x_{1}+x_{2}$
$X_{1}+x_{2}$ $\mathop{\rightarrow}\limits X_{1}$
$X_{2}+x_{1}$ $\mathop{\rightarrow}\limits X_{2}$
$x_{1}+x_{2}$ $\mathop{\rightarrow}\limits x_{1}$
It has bounded executions from every configuration: $\min(\#X_{1},\#X_{2})$ of the first reaction can occur, and the other reactions decrease molecular count, so are limited by the total configuration size. However, it is collapsing since, for any $n$, there exists an input of size $n$ that reaches a stable configuration of size 1. Theorem˜6.7 is similarly false without the all-voting hypothesis; for each of the reactions with one product above, add another non-voting product $W$. This converts the CRD to be non-collapsing but not all-voting. Of course, the execution bounded hypothesis is also necessary: the original population protocols paper [angluin2004computation] showed that all-voting, non-collapsing, leaderless population protocols can stably decide all semilinear predicates.
The following collapsing, all-voting, leaderless (but entirely execution bounded) CRD stably decides parity. Let the input species be named $X_{1}$. Species $X_{1}$ votes yes, $X_{0}$ votes no:
$X_{1}+X_{1}$ $\mathop{\rightarrow}\limits X_{0}$
$X_{1}+X_{0}$ $\mathop{\rightarrow}\limits X_{1}$
$X_{0}+X_{0}$ $\mathop{\rightarrow}\limits X_{0}$
It has bounded executions from any configuration: exactly $X_{1}+\#X_{0}-1$ reactions can occur since each reduces $X_{1}+\#X_{0}$ by 1. Similar to above, by adding the non-voting product $W$ to each reaction above, this CRD becomes non-collapsing but not all-voting, showing that the all-voting hypothesis is also necessary for stably deciding parity.
6.3 Impossibility of stably deciding not eventually constant predicates
We now present our main negative result, LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant, which generalizes Theorem˜6.7 to show that such CRNs can stably decide only very limited (eventually constant) predicates.
Definition 6.9.
Let $\phi:\mathbb{N}^{d}\rightarrow\{0,1\}$ be a predicate. We say $\phi$ is eventually constant if there is $n_{0}\in\mathbb{N}$ such that $\phi$ is constant on $\mathbb{N}_{\geq n_{0}}^{d}=\left\{\mathbf{x}\in\mathbb{N}^{d}\mid(\forall i\in\{1,\ldots,d\})\ \mathbf{x}(i)\geq n_{0}\right\}$, i.e., either $\phi^{-1}(0)\cap\mathbb{N}_{\geq n_{0}}^{d}=\emptyset$ or $\phi^{-1}(1)\cap\mathbb{N}_{\geq n_{0}}^{d}=\emptyset$.
In other words, although $\phi$ may have an infinite number of each output, “sufficiently far from the boundary of the positive orthant” (where all coordinates exceed $n_{0}$), only one output appears. See LABEL:fig:almost_vs_eventually_vs_not_constant for a 2D example.
Theorem 0.
thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant If a non-collapsing, all-voting, entirely execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
Proof sketch.
This proof is similar to that of Theorem˜6.7. In that proof, we repeatedly add a “constant amount of additional input $\{X_{2}\}$ or $\{X_{1}\}$, which flips the output”. For more general semilinear, but not eventually constant, predicates, we dig into the structure of the semilinear set to find a sequence of constant-size vectors representing additional inputs that flip the correct output. Any predicate that is not eventually constant has infinitely many yes inputs and infinitely many no inputs, but in general they could be increasingly far apart: e.g., $\phi(\mathbf{x})=1$ if and only if $2^{n}\leq\|\mathbf{x}\|<2^{n+1}$ for even $n$. For the potential function argument to work, each subsequent input needs to be at most a constant larger than the previous.
But if $\phi$ is semilinear (and not eventually constant) then we can show that there is a sequence of increasing inputs $\mathbf{x}_{0}\leq\mathbf{x}_{1}\leq\mathbf{x}_{2}\leq\dots$, each a constant distance from the next ($\|\mathbf{x}_{j+1}-\mathbf{x}_{j}\|=O(1)$), flipping the output ($\phi(\mathbf{x}_{j})\neq\phi(\mathbf{x}_{j+1})$). Roughly, this is true for one of two reasons. Using Theorem˜2.5, $\phi$ is a Boolean combination of threshold and mod sets. Either the mod sets are not combined to be trivially $\emptyset$ or $\mathbb{N}^{d}$, in which case we can find some vector $\mathbf{v}$ that, followed infinitely far from some starting point $\mathbf{x}_{0}$ (so $\mathbf{x}_{i}=\mathbf{x}_{0}+i\mathbf{v}$) periodically hits both yes inputs ($\phi(\mathbf{x}_{j})=1$) and no inputs ($\phi(\mathbf{x}_{j})=0$). (See LABEL:fig:periodic_set_2d and LABEL:fig:region_mod_periodic_not_constant_but_constant_along_recession_cone.) Otherwise, the mod sets can be removed and simplify the Boolean combination to only threshold sets, in which case the infinite sequence $\mathbf{x}_{0},\mathbf{x}_{1},\dots$ can be obtained by moving along a threshold hyperplane that separates yes from no inputs. (See LABEL:fig:threshold_sets.) ∎
Proof 6.11.
This proof is similar to that of Theorem˜6.7, with the vectors $\mathbf{v}_{i}$ defined below playing the role of the “constant amount of additional input $\{X_{2}\}$ or $\{X_{1}\}$ that flips the correct answer” in that proof.
Let $\mathcal{D}=(\Lambda,R,\Sigma,\Upsilon_{\mathrm{Y}},\Upsilon_{\mathrm{N}},\mathbf{s})$ be a CRD obeying the stated conditions, and suppose for the sake of contradiction that $\mathcal{D}$ stably decides a semilinear predicate $\phi$ that is not eventually constant.
By LABEL:lem:semilinear-not-almost-constant-alternating-period-vector-and-offset, there is an infinite sequence $\mathbf{x}_{0},\mathbf{x}_{1},\dots$ such that
  1. 1.
    $\phi(\mathbf{x}_{i})\neq\phi(\mathbf{x}_{i+1})$ (correct answer swaps for each subsequent input),
  2. 2.
    $\mathbf{x}_{i}\leq\mathbf{x}_{i+1}$, (inputs are increasing (on at least one coordinate(s)), and
  3. 3.
    for some constant $c$, $\|\mathbf{x}_{i+1}-\mathbf{x}_{i}\|\leq c$. (adjacent inputs are “close”)
Assume WLOG that $\phi(\mathbf{x}_{0})=0$. For each $i\in\mathbb{N}$, let $\mathbf{v}_{i}=\mathbf{x}_{i+1}-\mathbf{x}_{i}$, noting by condition (LABEL:lem:semilinear-not-almost-constant-alternating-period-vector-and-offset:cond2) that $\mathbf{v}_{i}\geq 0$.
We consider the sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\mathbf{a}_{2},\dots$ defined as follows. Let $\mathbf{a}_{0}$ be a stable configuration reachable from $\mathbf{x}_{0}$; since the correct answer is no, all species present in $\mathbf{a}_{0}$ vote no. Now add $\mathbf{v}_{0}$ to $\mathbf{a}_{0}$. By additivity, the configuration $\mathbf{a}_{0}+\mathbf{v}_{0}$ is reachable from $\mathbf{x}_{1}=\mathbf{x}_{0}+\mathbf{v}_{0}$. Since the correct answer for $\mathbf{x}_{1}$ is yes, $\mathcal{D}$ must go from $\mathbf{a}_{0}+\mathbf{v}_{0}$ to a stable “yes” configuration, call this $\mathbf{a}_{1}$. Now add $\mathbf{v}_{1}$ to $\mathbf{a}_{1}$. Since the correct answer is no, $\mathcal{D}$ must now reach from $\mathbf{a}_{1}+\mathbf{v}_{1}$ to a stable “no” configuration, call it $\mathbf{a}_{2}$. By condition (LABEL:lem:semilinear-not-almost-constant-alternating-period-vector-and-offset:cond3), each $\mathbf{v}_{i}$ obeys $\|\mathbf{v}_{i}\|<c$ for some constant $c$.
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\dots$ where all species in $\mathbf{a}_{i}$ vote yes for odd $i$, and all species in $\mathbf{a}_{i}$ vote no for even $i$. Since $\mathcal{D}$ is non-collapsing, the size of the configurations $\mathbf{a}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically nondecreasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.)
Since all species vote, for some constant $\delta>0$, to get from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$, at least $\delta\|\mathbf{a}_{i}\|$ reactions must occur. This is because all species in $\mathbf{a}_{i}$ must be removed since they vote the opposite of the voters in $\mathbf{a}_{i+1}$, and each reaction removes at most $O(1)$ molecules. (Concretely, let $\delta=1/\max_{(\mathbf{r},\mathbf{p})\in R}\|\mathbf{r}\|-\|\mathbf{p}\|$, i.e., 1 over the most net molecules consumed in any reaction.)
Since $\mathcal{D}$ is entirely execution bounded, by LABEL:thm:potential-function-iff-entirely-execution-bounded, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{w}\cdot\mathbf{x}$, where $\mathbf{w}\geq\mathbf{0}$. Adding $\mathbf{v}_{i}$ to $\mathbf{a}_{i}$ increases $\Phi$ by $\mathbf{w}(\mathbf{v}_{i})$, which is bounded above by a constant since $\|\mathbf{v}_{i}\|<c$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $\mathbf{v}_{i}$ and then going from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{w}(\mathbf{v}_{i})+2$).
However, $\Phi$ started at the constant $\Phi(\mathbf{x}_{0})$. Before $\|\mathbf{a}_{i}\|$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{w}(\mathbf{v}_{i})+2$ (i.e., large enough that the net change in $\Phi$ is negative resulting from adding a single input and going to the next stable configuration), $\Phi$ could increase, if $\Phi(\mathbf{v}_{i})$ is larger than the net decrease in $\Phi$ due to following reactions to get from $\mathbf{a}_{i}+\mathbf{v}_{i}$ to $\mathbf{a}_{i+1}$.
However, since $\mathcal{D}$ is non-collapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{x}_{0})$), after which point $\Phi$ strictly decreases after each round of this process.
At some point in this process, $\mathcal{D}$ will not be able to reach all the way to the next $\mathbf{a}_{i}$ without $\Phi$ becoming negative, a contradiction.
The statement of Theorem˜6.7 does not mention the concept of a leader, but it would typically apply to leaderless CRDs. A CRD may be execution bounded from configurations with a single leader, but not execution bounded when multiple leaders are present (preventing the use of LABEL:thm:potential-function-iff-entirely-execution-bounded, which requires the CRD to be execution bounded from all configurations). For example, in Lemma˜4.7, reaction (9) occurs finitely many times if the leader/voter $S_{Y}$ or $S_{N}$ has count 1. However, if $S_{Y}$ and $S_{N}$ can be present simultaneously (e.g., if we start with two leaders), then the reactions $S_{Y}+V_{NN}\mathop{\rightarrow}\limits S_{Y}+V_{YN}$ and $S_{N}+V_{YN}\mathop{\rightarrow}\limits S_{N}+V_{NN}$ can flip between $V_{NN}$ and $V_{YN}$ infinitely often in an unbounded execution.
If the CRN is leaderless, however, we have the following, which says that if it is execution bounded from valid initial configurations, then it is execution bounded from all configurations.
Lemma 6.12.
lem:leaderless-and-execution-bdd-implies-entirely-execution-bdd If a leaderless CRD or CRC is execution bounded, then it is entirely execution bounded.
Proof sketch.
Since $\mathcal{C}$ is leaderless, the sum of two valid initial configurations is also valid. Thus if we can produce some species from a valid initial configuration, we can produce arbitrarily large counts of all species by adding up sufficiently many initial configurations. This means that for any configuration $\mathbf{x}$, from any sufficiently large valid initial configuration $\mathbf{i}$, some $\mathbf{y}\geqq\mathbf{x}$ is reachable from $\mathbf{i}$. But if $\mathcal{C}$ is execution bounded from $\mathbf{i}$, since $\mathbf{i}\Rightarrow\mathbf{y}$, it must also be execution bounded from $\mathbf{y}$, thus also from $\mathbf{x}$ since by additivity any reactions applicable to $\mathbf{x}$ are also applicable to $\mathbf{y}$. ∎
Proof 6.13.
Let $\mathcal{C}$ be a leaderless CRD or CRC. Let $\mathbf{x}$ be any configuration. We first show that some $\mathbf{y}\geqq\mathbf{x}$ is reachable from a valid initial configuration $\mathbf{i}$.
We may assume without loss of generality that $\mathcal{C}$ only contains species producible from valid initial configurations, otherwise we obtain an equivalent CRN by removing those unproducible species from $\mathcal{C}$.
Since $\mathcal{C}$ is leaderless, the sum of two valid initial configurations is also valid. Then each species $S$ being producible means that there is a valid initial configuration $\mathbf{i}_{S,1}$ such that for some $\mathbf{y}_{S,1}$, $\mathbf{i}_{S,1}\Rightarrow\mathbf{y}_{S,1}$ and $\mathbf{y}_{S,1}(S)\geq 1$, i.e., at least one copy of $S$ can be produced. Let $\mathbf{i}_{S,k}=k\cdot\mathbf{i}_{S,1}$. By additivity, $\mathbf{i}_{S,k}\Rightarrow\mathbf{y}_{S,k}$, where $\mathbf{y}_{S,k}=k\cdot\mathbf{y}_{S,1}$, noting that $\mathbf{y}_{S,k}(S)\geq k$. In other words, all species are producible in arbitrarily large counts from some valid initial configuration.
Now we argue all species can be made simultaneously arbitrarily large count from some valid initial configuration; in particular, we can reach a configuration with counts at least $\mathbf{x}$. Let $\mathbf{i}=\sum_{S\in\Lambda}\mathbf{i}_{S,\mathbf{x}(S)}$. Since each $\mathbf{i}_{S,\mathbf{x}(S)}\Rightarrow\mathbf{y}_{S,\mathbf{x}(S)}$, by additivity we have $\mathbf{i}\Rightarrow\mathbf{y}$, where $\mathbf{y}=\sum_{S\in\Lambda}\mathbf{y}_{S,\mathbf{x}(S)}$. Then for each $S\in\Lambda$, $\mathbf{y}(S)\geq\mathbf{x}(S)$, so $\mathbf{y}\geqq\mathbf{x}$.
Since all executions from $\mathbf{i}$ are finite, all executions from $\mathbf{y}$ are finite. By additivity, any sequence of reactions applicable to $\mathbf{x}$ is also applicable to $\mathbf{y}$. Thus all executions from $\mathbf{x}\leqq\mathbf{y}$ must be finite as well, i.e., $\mathcal{C}$ is entirely execution bounded since $\mathbf{x}$ is an arbitrary configuration.
LABEL:lem:leaderless-and-execution-bdd-implies-entirely-execution-bdd lets us replace “entirely execution bounded” in LABEL:thm:non-collapsing_allvoting_entirely_execution_bounded_decide_almost_constant with “leaderless and execution bounded”:
Corollary 6.14.
If a non-collapsing, all-voting, leaderless, execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
In particular, since the original model of population protocols [angluin2004computation] defined them as leaderless and all-voting—and since population protocols are non-collapsing—we have the following.
Corollary 6.15.
If an execution bounded population protocol stably decides a predicate $\phi$, then $\phi$ is eventually constant.
7 Conclusion
We explored the computational capabilities of execution bounded Chemical Reaction Networks (CRNs), which terminate after a finite number of reactions. This constraint aligns the model with practical scenarios where fuel supply is limited.
Our findings illustrate that the computational power of these CRNs varies significantly based on structural choices. Specifically, CRNs with an initial leader and the ability to allow only the leader to vote can stably compute all semilinear predicates and functions in $O(\|x\|\log\|x\|)$ parallel time. Without an initial leader, and requiring all species to vote, these networks are limited to computing eventually constant predicates. This limitation holds considerable weight for decentralized systems modeled by population protocols, which inherently exhibit these traits. Additionally, we introduced a new characterization of execution bounded networks through a nonnegative linear potential function, providing a novel theoretical tool for analyzing the physical constraints CRNs.
A key question remains open: Can execution bounded CRNs compute semilinear functions and predicates within polylogarithmic time? Angluin, Aspnes and Eisenstat [AngluinAE2008Fast] introduced a fast population protocol that simulates a register machine with high probability, and can be made probability 1 with semilinear predicates. However, this construction seems inherently unbounded in executions.